OpenStudy (roxanne):

complete the table of ordered pairs for each linear equation x=-y+4 let y=0

jhonyy9 (jhonyy9):

x=4

OpenStudy (anonymous):

x=-y+4 x=-(0)+4 x=4

OpenStudy (anonymous):

x=4